# A defect-chemistry-informed phase-field model of grain growth in oxide electroceramics

### v2

@@ Abstract
Dopants can significantly affect the properties of oxide ceramics through their impact on the property-determined microstructure characteristics such as grain boundary segregation, space charge layer formation in the grain boundary vicinity, and the resultant microstructure features like bimodality due to abnormal grain growth. To support rational oxide ceramics design, we propose a multiphysics-based and defect-chemistry-informed phase-field grain growth model to simulate the microstructure evolution of oxide ceramics. It fully respects the thermodynamics of charged point defects (oxygen vacancies and dopants) in both the grain interior and boundaries and considers the competing kinetics of defect diffusion and grain boundary movement. The proposed phase-field model is benchmarked against well-known simplified bicrystal models, including the Mott-Schottky and Gouy-Chapman models. Various simulation results are presented to reveal the impacts of defect formation energy differences between the grain interior and the grain boundary core on the key microstructural aspects. In particular, simulation results confirm that the solute drag effect alone can lead to bimodal grain size distribution, without any contribution from grain misorientation and other anisotropy. Interestingly, abnormal grain growth simulations demonstrate that grain boundary potentials can vary substantially: grain boundaries of larger grains tend to have lower potentials than those of smaller grains. Such heterogeneous grain boundary potential distribution may inspire a new material optimization strategy through microstructure design. This study provides a comprehensive framework for defect-chemistry-consistent investigations of microstructure evolution in polycrystalline oxide ceramics, offering fundamental insights into in-situ processes during critical manufacturing stages.
@@
Phase-field method , Defect-chemistry , Space charge layer , Abnormal grain growth
@@ 1 Introduction
In recent decades, oxide ceramics, such as SrTiO3 (STO) and BaTiO3 (BTO), have undergone tremendous development and are increasingly crucial in applications, such as solid oxide fuel cells [1, 2], sensors [3, 4], capacitors [5, 6], actuators [7, 8], and memristors [9, 10]. Oxide ceramics can be fabricated via sintering or deposition processes [11, 12, 13, 14]. When manufactured through sintering, oxide ceramics exhibit a polycrystalline microstructure, with their macroscopic properties heavily influenced by the synthesis conditions [15]. To tailor the physical and chemical properties of oxide ceramics, heterovalent substitution (doping) is an effective method for introducing additional charged defects through ionic compensation during synthesis. The state-of-the-art designs are mostly empirical, since knowledge of the correlation among doping strategies, processing and the resultant microstructure is limited. The influential mechanisms of doping on microstructure in oxide ceramics are manifold. First of all, the addition of doping elements in the parent material systems triggers defect chemistry reaction and results in multiple types of charged point defects or defect complexes, which interact with the grain boundary (GB) chemically, electrically and elastically and form space charge layers (SCLs) around GB. More specifically, the equilibrium concentrations of the point defects within the bulk are thermodynamically determined through defect chemistry [16, 17]. The scenario near inside the GB core and in the GB vicinity are different. In undoped and acceptor doped perovskites such as Fe-doped STO, the GB core is positively charged with excess oxygen vacancies, notated $\text{V}_{\text{O}}^{\text{..}}$ in Kröger-Vink notation [18], stemming from the negative formation energy difference of oxygen vacancy between GB core and bulk, i.e. $\Delta g_{\text{V}_{\text{O}}^{..}}=\mu^{0}_{\text{V}_{\text{O}}^{..},\text{c}}-\mu^{0}_{\text{V}_{\text{O}}^{..},\text{b}}<0$. To compensate the positively charged GB core, a negatively charged region forms adjacent to it, called the space charge layer (SCL).
@@
Sharp interface models have been used to describe the SCL and the related grain boundary potential. Among them are the Mott-Schottky (MS) model and the Gouy-Chapman (GC) model. The MS model assumes that the dopant concentration is constant, and only the oxygen vacancies are able to redistribute, in order to form the SCLs. In contrast, the GC model has both defects as mobile and thus allows both to contribute to charge compensation. Unlike oxygen vacancies, acceptor dopant forms an accumulation region within the SCL and in the GB core due to their negative charge [Fig. 1(e)]. The spatial inhomogeneity introduced by SCL formation exerts a profound impact on the transport dynamics of charge species. Evidence shows that cation diffusion parallel to the GB is expedited [19, 20], whereas charge transport across the GB is significantly hindered [21].
@@
SCLs determine not only the GB local properties of the synthesized material, but also impact the grain growth dynamics already during the synthesis, e.g., the well-known solute drag effect. Grain growth in oxide ceramics, driven by a reduction in surface area and controlled by the diffusion of charged species, plays a crucial role in determining the final grain size distribution and the concentration distributions of various point defects. Segregation in a moving GB core necessitates additional free energy dissipation and results in the solute drag effect [22, 23]. This effect significantly interferes with the kinetics of grain growth and can independently trigger abnormal grain growth even in the absence of any texture or pinning sites [24]. Dopant segregation to GB core has been experimentally evidenced, especially at high temperatures [25, 26]. Experiments also show that even 0.2% amount of Fe in STO can pin certain portion of GBs [Fig. 1(b)]. This trend increases with 2% Fe [Fig. 1(c)]. When increasing the dopant concentration to 5% [Fig. 1(d)], the grain growth is completely suppressed [25, 27]. In addition, the grain growth coefficient is found to follow a non-Arrhenius behavior between $1350^{\circ}$ and $1425^{\circ}$. Two types of GB core with different mobilities are proposed to explain the non-Arrhenius behavior [28, 25]. Despite various evidence supporting the impact of SCLs on GB properties and grain growth dynamics, experimental characterization of the SCL remains a challenge [29]. The complexity arises from the inherent structural and compositional heterogeneities present in polycrystalline oxide ceramics. One of the primary obstacles in this characterization is the limitation of in-situ transmission electron microscopy (TEM) investigations. Such studies are typically unfeasible due to the high sintering temperatures required for these ceramics, which exceed the operational limits of most TEM setups.
It is hence desirable to employ theoretical models and numerical simulations to evaluate SCLs and abnormal grain growth. De Souza derived numerical solutions based on defect chemistry in acceptor-doped oxide systems at equilibrium state [30]. The impacts of different charge species, the formation energy difference of oxygen vacancy and dopant between GB core and bulk phase, temperature and partial pressure of oxygen on grain boundary potential are comprehensively studied in a bicrystal with planar GB core. However, applying these numerical solutions to polycrystalline cases is impractical. Under this circumstance, the phase-field method with a diffuse interface emerges as a powerful tool to solve different moving boundary problems [31, 32]. There are a series of phase-field sintering models, e.g. to consider the solute drag effect in metals [24] and ceramics [33] and for non-isothermal sintering [34, 35, 36, 37]. But these models do not consider the charged defects, electrostatics or defect chemistry, which are important for oxide ceramics. Recently, Vikrant et al. developed a multi-physics phase-field model with an emphasis on GB misorientation and investigate the SCL formation at equilibrium and quasi-equilibrium state [38]. Then, this model is applied to explain the abnormal grain growth with a bimodal pattern as observed in experiments [39]. However, in this phase-field model the defect chemistry is not explicitly included, and thus there is no simulation results dedicated to the influence of defect energetics on SCLs and abnormal grain growth.
@@
In the present work, we develop a phase-field grain growth model founded on the principle of defect-chemistry to simulate SCL formation and abnormal grain growth during sintering of oxide electroceramics. By taking into account the formation energy differences and the equilibrium defect concentrations in bulk and GB core, the thermodynamic descriptions of point defects are included in the phase-field model, along with the kinetics of grain boundary motion and defect diffusion. The spatio-temporal distributions of defect concentration and electrostatic potential in the grain interior and at GB are then solved self-consistently.
The outline of this paper is as follows. Section 2 presents the fundamentals of defect chemistry and the classical sharp interface bicrystal models, i.e., the GC and MS models, for they serve as important benchmarks in later sections. Section 3 describes the grain growth phase-field model in the principle of defect chemistry and electrostatics. Then, we demonstrate the implementation of the phase-field model via the finite element method in Section 4. The capabilities of this model are discussed for bicrystal and polycrystalline cases in Section 5. Thereby the solute drag induced abnormal grain growth simulation results are systematically presented and discussed, particularly in terms of the impact of segregation energies. It is followed by a concluding Section 6.

### v3

# A defect-chemistry-informed phase-field model of grain growth in oxide ceramics: application to Fe-doped SrTiO3

@@ Abstract
Dopants can significantly affect the properties of oxide ceramics through their impact on the property-determined microstructure characteristics such as grain boundary (GB) segregation, space charge layer formation in the GB vicinity, and the grain growth deviating from normal patterns. To support the rational design of oxide ceramics, we propose a defect-chemistry-informed phase-field grain growth model to simulate the microstructure evolution of oxide ceramics. It fully respects the defect-chemistry theory by accounting for the distinct segregation energies and available site densities of charged point defects (oxygen vacancies and acceptor dopants) in both the grain interior and boundaries, and it considers the competing kinetics of defect diffusion and GB movement. The proposed phase-field model is benchmarked against well-known bicrystal models, including the Mott-Schottky and Gouy-Chapman models. Various simulation results are presented to reveal the effect of different defect-chemistry parameters on the space charge layer formation and key microstructural aspects. In particular, simulation results confirm that the solute drag effect alone can lead to skewed grain size distribution that do not follow the log-normal distribution, without any contribution from grain misorientation and other anisotropy. Interestingly, simulations also demonstrate that grain boundary potentials can vary substantially: grain boundaries of larger grains tend to have lower potentials than those of smaller grains. Such heterogeneous grain boundary potential distribution may inspire a new material optimization strategy through microstructure design. This study provides a comprehensive framework for defect-chemistry-consistent investigations of microstructure evolution in polycrystalline oxide ceramics, offering fundamental insights into microscopic processes during critical manufacturing stages.
@@
Phase-field method , Defect-chemistry , Space charge layer , Grain growth
@@ 1 Introduction
In recent decades, oxide ceramics, such as SrTiO3 (STO) and BaTiO3 (BTO), have undergone tremendous development and are increasingly crucial in applications, such as capacitors lee2011atomic , hou2017ultrahigh , actuators kursumovic2013new , baek2013epitaxial , sensors wang2022stretchable , chan2014highly , memristors muenstermann2010coexistence , hu2013ferroelectric and electrolytes in fuel cells ishihara1994doped , kreuer2003proton , li2014family , Oxide ceramics can be fabricated via sintering or deposition mishra2022ultra , reavley2022ultrafast , sayer1990ceramic , li2003two . When manufactured through sintering, oxide ceramics exhibit a polycrystalline microstructure, with their macroscopic properties heavily influenced by the synthesis conditions chaim2008sintering . To tailor the physical and chemical properties of oxide ceramics, heterovalent substitution (doping) is an effective method for introducing additional charged defects through ionic or electronic compensation during synthesis. The state-of-the-art designs are mostly empirical, since knowledge of the relationships between doping strategies, processing and the resultant microstructure is limited. The influential mechanisms of doping on microstructure in oxide ceramics are manifold. First of all, the addition of doping elements in the parent material systems triggers defect chemistry reactions and results in multiple types of charged point defects or defect complexes, which interact with the grain boundary (GB) chemically, electrically and elastically and form space charge layers (SCLs) around GB. More specifically, the equilibrium concentrations of the point defects within the bulk are thermodynamically determined through defect chemistry maier2023physical , smyth2000defect . The scenario near inside the GB core and in the GB vicinity are different. In undoped and acceptor doped perovskites such as Fe-doped STO, the GB core is positively charged with excess oxygen vacancies, $\text{V}_{\text{O}}^{\text{..}}$ in Kröger-Vink notation kroger1956relations , stemming from the negative formation energy difference of oxygen vacancy between GB core and bulk, i.e. $\Delta g_{\text{V}_{\text{O}}^{..}}=\mu^{0}_{\text{V}_{\text{O}}^{..},\text{c}}-\mu^{0}_{\text{V}_{\text{O}}^{..},\text{b}}<0$. To compensate the positively charged GB core, negatively charged regions form adjacent to it, called the space charge layers (SCLs).
@@
Sharp interface models have been used to describe the SCL and the related grain boundary potential. Among them are the Mott-Schottky (MS) model mott1939theory , mott1938note , schottky1939halbleitertheorie and the Gouy-Chapman (GC) model guoy1910constitution , chapman1913li . The MS model assumes that the dopant concentration is constant, and only the oxygen vacancies are able to redistribute, in order to form the SCLs. In contrast, the GC model has both defects as mobile and thus allows both to contribute to charge compensation. While positively charged oxygen vacancies are depleted in the case of a positive core, acceptor dopants are depleted in the case of a positive core within the SCL and in the GB core due to their negative charge [Fig. 1(e)]. The spatial inhomogeneity introduced by SCL formation exerts a profound impact on the transport dynamics of charge species. Evidence shows that cation diffusion parallel to the GB is expedited parras2020grain , savzinas2017134ba , whereas charge transport across the GB is significantly hindered waser2000grain .
@@
SCLs determine not only the GB local properties of the synthesized material, but also impact the grain growth dynamics already during the synthesis through the well-known solute drag effect cahn1962impurity . Grain growth in oxide ceramics, driven by a reduction in surface area and controlled by the diffusion of charged species, plays a crucial role in determining the final grain size distribution and the concentration distributions of various point defects. Segregation in a moving GB core necessitates additional free energy dissipation and results in the solute drag effect cahn1962impurity , hillert1976treatment . This effect strongly affects the kinetics of grain growth and can independently trigger abnormal grain growth even in the absence of any texture or pinning sites kim2008grain . Dopant segregation to GB core has been experimentally evidenced, especially at high temperatures zahler2023grain , jennings2024grain . Experiments also show that even 0.2% amount of Fe in STO can pin a certain portion of GBs [Fig. 1(b)]. This trend increases at 2% Fe [Fig. 1(c)]. When increasing the dopant concentration to 5% [Fig. 1(d)], the grain growth is completely suppressed zahler2023grain , rheinheimer2016grain . In addition, the grain growth coefficient is found to follow a non-Arrhenius behavior between $1350^{\circ}$C and $1425^{\circ}$C. Two types of GB core with different mobilities have been proposed to explain the non-Arrhenius behavior rheinheimer2015non , zahler2023grain . Despite various pieces of evidence supporting the impact of SCLs on GB properties and grain growth dynamics, experimental characterization of the SCL remains a challenge fleig2002grain . The complexity arises from the inherent structural and compositional heterogeneities present in polycrystalline oxide ceramics. One of the primary obstacles in this characterization is the limitation of in-situ transmission electron microscopy (TEM) investigations. Such studies are typically unfeasible due to the high sintering temperatures required for these ceramics, which exceed the operational limits of most TEM setups.
It is hence desirable to employ theoretical models and numerical simulations to evaluate SCLs and abnormal grain growth. Kliewer and Koehler derived the local density of charge carriers in the presence of electrostatic potentials from a global approach kliewer1965space . Through minimizing the free energy of the entire crystal, the equations for the densities of various defects are derived in the type of Boltzmann distribution. Blakely and Macdonald extended the global thermodynamic approach with the consideration of an exhaustible number of defect sites and derived Fermi-Dirac-like distribution functions of defects blakely1973space , mukhopadhyay1991ionic , macdonald1980interfacial . Jamnik employed a local thermodynamic formalism, in which the interface region was divided into a core and a space charge region with an assumption that the material constants, i.e. the standard chemical potentials and mobilities of defects, behave as step functions jamnik1995interfaces . De Souza extended the local thermodynamic approach to abrupt GB||SCL model based on defect chemistry in acceptor-doped oxide systems at equilibrium state de2009formation . The impacts of different charge species, the formation energy difference of oxygen vacancy and dopant between GB core and bulk phase, the number of available sites of oxygen vacancy and dopants, temperature and partial pressure of oxygen on grain boundary potential are comprehensively studied in a STO bicrystal with planar GB core de2009formation . Then, the abrupt GB||SCL model was implemented to reproduce the SCL in STO to investigate the impact of negative electrode on the defects distribution, GB potential and grain growth kinetics rheinheimer2019grain . In addition, this model was also employed to investigate the SCL formation in fluorite-type oxide parras2020grainJACS , tong2020analyzing , parras2020grain .
However, applying the abrupt GB||SCL model to polycrystalline sintering scenarios is challenging. The complex microstructure, intricate GB distribution and GB kinetics affected by solute drag effects hinder the practical use of the sharp interface methods. Under this circumstance, the phase-field method with a diffuse interface emerges as a powerful tool to solve different moving boundary problems wang2020modeling , wang2021quantitative , hu2024data . Kobayashi et al. developed the Kobayashi-Warren-Carter (KWC) phase-field model to consider crystallographic orientation during grain growth kobayashi2000continuum . Moelans et al. proposed a homogeneous multi-well potential in the free energy density functional to obtain the quantitative relations between phase-field model parameters and GB properties moelans2008quantitative . Additionally, there are also a series of phase-field sintering models for non-isothermal sintering yang20193d , yang2020investigation , yang2022diffuse , oyedeji2023variational . Several researchers have developed phase-field models to capture solute drag effects during grain growth processes. The first phase-field simulation was performed by Fan et al. fan1999computer . Their model revealed that solute drag introduces a nonlinear relationship between GB velocity and driving force and influences the growth exponent in polycrystalline systems. Cha et al. proposed a phase-field model that treats the GB as a distinct phase and studied the impact of solute drag on migrating grain boundaries in a binary alloy system cha2002phase . The effect of solute drag is automatically incorporated into the model and it is capable of reproducing the equilibrium solute segregation and the free energy dissipation. Grönhagen et al. introduced a concentration dependency in the height of the double-well potential in the Gibbs-energy expression and verified the possibility of modeling solute drag on a moving boundary gronhagen2007grain . This model was then adopted by Kim kim2008grain and Li li2009phase . Kim et. al investigated the grain growth in association with GB segregation in two-dimensional polycrystalline systems kim2008grain . They concluded abnormal grain growth can be triggered by the solute drag effect. Li et al. observed that the solute concentration at the moving GB may increase with increasing velocity and becomes larger than the equilibrium value, which is unexpectedly predicted by solute drag theory li2009phase . They also found non-linear relationship between the curvature driving force and the GB velocity can exist not only in the transition regime of the velocity, but also in the low velocity regime. Greenwood et al. studied the atomic scale interaction of solute with a migrating GB by using a phase-field crystal model greenwood2012phase . They developed a new formalism to allow for the application of an external driving pressure for the growth of one grain. Extending from grain growth studies, phase-field models for solid-state sintering in the initial and final stage considering various diffusion paths are proposed by Wang et al. wang2006computer , Kumar et al. kumar2010phase , Hötzer et al. hotzer2019phase and Rehn et al. rehn2019phase . Rheinheimer et. al employed a multi-phase-field model to simulate the abnormal grain growth in STO ceramics by introducing two types of GB mobilities rheinheimer2019non . However, the electrostatic effects related SCL is not considered.
In order to consider the electrostatic contribution in the phase-field model, Guyer et al. developed a phase-field model to capture the charge distribution in an equilibrium double layer at the electrochemical interface with mass and volume constraints, Poisson’s equation, ideal solution thermodynamics, and a simple description of the competing energies in the interface guyer2004equilibrium . By using the same model, Guyer et al. also explored its kinetic behavior for electrodeposition and electrodissolution conditions guyer2004kinetic . They demonstrated the ohmic conduction in the electrode and ionic conduction in the electrolyte and found the nonlinear relationship between current and overpotential as predicted by Butler-Volmer equation. Bishop et al. formulated a Helmholtz free energy functional with the electrostatic contribution and studied the phase separation of charged species by spinodal decomposition bishop2003effect . García et al. incorporated the laws of thermodynamics and Maxwell’s equations and proposed a theoretical framework including thermodynamically consistent equilibrium equations and kinetic driving forces to describe the time evolution for electrically and magnetically active materials garcia2004thermodynamically . Lund et al. proposed a generalized framework that integrates the free energy contributions of thermochemical, structural, mechanical, and electrical fields lund2021thermodynamically . They calculated the SCL in gadolinium-doped cerium oxide with high dopant concentration and predicted the macroscopic ionic conductivity with different segregation energy of oxygen vacancy.
Aiming to investigate the sintering process with electrostatics, multi-physics phase-field model needs to be developed. Vikrant et al. developed a phase-field model including the electro-chemo-mechanical effects to physically describe the equilibrium and transport properties of charged interfaces in ion-conducting solids vikrant2018charged . They simulated gadolinium-doped cerium oxide and investigated the influence of gadolinia substitution concentration on the SCL. Then, Vikrant et al. extended this model to understand the effects of the intrinsic and extrinsic ionic species and point defects on the structural and electrochemical stability of grain boundaries in polycrystalline ceramics vikrant2019charged . They studied the impacts of different crystallographic misorientations on SCL in the cubic Yttria Stabilized Zirconia. Recently, Vikrant et al. developed a phase-field model based on KWC grain growth model and investigate the SCL formation at equilibrium and quasi-equilibrium state with the consideration of the GB misorientation and solute drag effect vikrant2020electrochemical . They considered the negative segregation formation energy of oxygen vacancies for SCL formation (and unfortunately treated oxygen vacancies and cation dopants as sharing the same crystallographic site). Then, this model was applied to explain the abnormal grain growth with a bimodal pattern as observed in experiments vikrant2020electrochemicallyActa . In addition to free energy based phase-field model, Aagesen et al. developed a grand potential-based phase-field model and included includes the effects of charged vacancies and the associated interactions between internal and applied electric fields aagesen2024electrochemical . In oxide ceramics such as Fe-doped STO, Fe atoms substitute for Ti4+ ions on the B-site of the perovskite sublattice, while charge compensation typically occurs via the formation of oxygen vacanies on the anion site. As a result, in the bulk phase, the number of available sites for an acceptor dopant corresponds to the number density of B-sites, which is $1/a^{3}$, where $a$ is the lattice parameter. The number of available sites indicates the maximum number of sites occupied by defects. In contrast, the number of available sites for an oxygen vacancy is three times higher, i.e. $3/a^{3}$. Even more important, B-site dopant and oxygen vacancy belong to different sub-lattices and thus have to be accounted for separately, in order to correctly represent the respective configurational entropy maier2023physical . In addition, the local structure within the GB core differs from that of the bulk. These structural differences can significantly alter the number of available sites for both dopants and oxygen vacancy. Therefore, it is necessary to treat the number of available sites for oxygen vacancies and dopants in the GB core as distinct from those in the bulk and these quantities play a critical role in determining the SCL. Variations in the number of available sites for oxygen vacancy and dopant give rise to diverse SCL formation behaviors de2009formation . Although Vikrant and Aagesen performed landmark studies on the space charge layer formation and grain microstructure evolution during the sintering process, the accurate coupling of these defect chemistry parameters into phase-field simulations was not addressed.
@@
Therefore, defect-chemistry consistent SCLs are characterized by a number of inequivalent defects, each with its unique segregation energy and number of available sites. Some GB cores exhibit not only a negative segregation formation energy for oxygen vacancies ($\Delta_{\text{seg}}g_{\text{V}_{\text{O}}^{..}}<0$), but also a combined negative segregation formation energy for oxygen vacancies and acceptor cations ($\Delta_{\text{seg}}g_{\text{V}_{\text{O}}^{..}}<0$ and $\Delta_{\text{seg}}g_{\text{dop}}<0$) usler2024point . Meanwhile, different defects in the bulk and GB core require different numbers of available sites usler2024point , usler2024space , de2019effect . Moreover, the dominant effect leading to abnormal grain growth during the sintering process in oxide electroceramics remains debatable. The combined effects of GB misorientation and solute drag have been explored in Fe-doped STO vikrant2020electrochemical , vikrant2020electrochemicallyActa . Nevertheless, an investigation of GB energy anisotropy by the GB grooving technique indicates that the reduction of GB energy due to anisotropy can only lead to moderate (about 15%) changes in grain growth rate kelly2018anti . Another investigation of GB mobility anisotropy pointed out the GB mobility anisotropy transition temperature lies in the abnormal grain growth regime and can impact the fast-type and low-type GBs. However, it does not seem as the grain growth transition is caused by a change in the GB anisotropy. Instead, studies suggest that the transition is associated with a change in the GB stoichiometry, while the atomistic GB structure remains largely unchanged through the grain growth transition sternlicht2016mechanism , shih2010investigation , baurer2010abnormal , sternlicht2019characterization , sternlicht2019characterization2 . Additionally, very recent experimental observations demonstrate that the abnormal grain growth in Fe-doped STO during sintering is strongly influenced by dopant concentration zahler2023grain . Nominally undoped STO, exhibits a monomodal grain size distribution despite the presence of GB misorientation. Hence, this concentration-dependent grain size distribution highlights the critical role of solute drag in driving abnormal grain growth, rather than GB misorientation.
Accordingly, we develop a phase-field grain growth model grounded in the principles of defect chemistry to reproduce SCL formation and investigate the role of solute drag effects during the sintering process. We exclude additional influences, such as GB misorientation on the grain growth process, on purpose, and in this way, we demonstrate that, even in the absence of these factors, the essential features of skewed grain size distribution which do not follow the normal grain growth can still be observed solely through the solute drag effect. The novelty of this work lies in three key aspects. First, this phase-field model accounts for unique segregation energies and distinct site densities of different defects in both the bulk phase and GB core, ensuring that the defect chemistry of both bulk and core are treated in a thermodynamically consistent description. Second, the capabilities of the phase-field model are rigorously benchmarked against sharp interface calculations, providing validation and accuracy. Third, the phase-field results reveal that solute drag alone is sufficient to lead to the skewed grain size distribution, confirming its role as the essential mechanism underlying this phenomenon.
The outline of this paper is as follows. Section 2 presents the fundamentals of defect chemistry and the classical sharp interface bicrystal models, i.e., the GC and MS models, for they serve as important benchmarks in later sections. Section 3 describes the grain growth phase-field model in the principle of defect chemistry and electrostatics. Then, we demonstrate the implementation of the phase-field model via the finite element method in Section 4. The capabilities of this model are discussed for bicrystal and polycrystalline cases in Section 5. Thereby the solute drag induced skewed grain size distribution are systematically presented and discussed, particularly in terms of the impact of different defect chemistry parameters. It is followed by a concluding Section 6.